# Robust Capacity Expansion Modelling for Renewable Energy Systems under Weather Uncertainty

### v3

@@ Methods
If an energy system model displays supply gaps during feasibility testing, its CAPEX optimisation problem is modified to obtain a robust solution. Consider a solution $\delta$ to $COMP^{i}_{j}$. In this solution, $1^{\intercal}\delta_{i}^{x}$ units of energy are missing. This information needs to be integrated into the original CAPEX problem to enable the energy system model to provide the extra energy supply required to meet demand.
Multiple modifications were iteratively developed to capture different sources of uncertainty. Initial testing showed a subset of them to be ineffective. In order to also report null–results, these modifications are outlined in Appendix E. The three promising modifications are inspired by three physical phenomena relevant to energy systems: 1) The importance of peak demands for energy system design 35, 36, 2) The existence of extended critical time periods 12, 13, 15, 14 and 3) variations in total yearly energy supply across different weather years 9.
For MOD 1 / Demand increase, update the demand vector $b_{j}$ of the starting weather year $j$ by adding the demand-supply gap $\delta$ to it, i.e.
@@
This adds the missing energy supply exactly when it is needed. One observation during initial testing was that this can lead to very large peaks in artificial demand, especially if done repeatedly over multiple iterations, leading to excessive battery installation. To counteract this, the artificial, additional demand can be divided up between neighbouring time periods, i.e. update $b$ and smooth $\delta$, i.e. $b^{\prime}=b+f(\delta)$ for some function $f:\mathbb{R}^{n}_{\geq 0}\mapsto\mathbb{R}^{n}_{\geq 0}\setminus\emptyset$ that changes $b$ locally. Figure 13 in Appendix E illustrates the effects of smoothing. Throughout our work, we differentiate between smoothed and non–smoothed MOD 1 / Demand increase.
Since feasibility tests show that supply gaps usually occur in adjacent time steps, those are clustered to identify time periods with potential supply gaps, i.e. critical time periods. Critical time periods are usually characterised by a combination of low electricity generation from renewables together with increased electricity demand, so–called (cold) dark lulls. A hierarchical clustering was chosen, as this has been shown to be effective for time–series aggregation previously 37, 38, and it allows for easy visual inspection and clustering along adjacent time–steps 39, 40. For each time period, check whether the optimised capacities provide sufficient energy. For this, calculate the potential electricity generation of the selected energy system model and compare this with the total demand in the time period. [^3] Note that a positive average supply gap means that the investigated energy systems lack dispatchable capacity, and thus the feasibility testing of the respective time periods will always find a supply gap. However, a time period with negative average hourly supply gaps can still contain supply gaps that are only found during feasibility testing. This happens if either total annual electricity generation is too low to cover overall demand, or if there are some demand peaks that can not be covered. For calculating the potential electricity generation, full load is assumed for generation and backup capacities. Let $b_{t,j}$ be the demand at hour $t\in T$, and $a_{t,j}^{k}\cdot x_{s}$ the time–series data and the calculated capacities of the energy system component $p\in P$ in weather year $u_{j}$, respectively. $P_{sup,con}=P_{sup}\cup P_{con}$ is used to denote that summation only happens over supply and conversion technologies. Then, for time interval $T\subseteq\mathcal{T}$, the average hourly supply gap $\Delta(T)_{j}$ for some weather year $u_{j}^{\prime}$ is given via
@@
for some critical time period $T$ and weather years $j,j^{\prime}$.
@@
If this does not suffice to achieve feasibility for some weather year $j$, we model long-term energy storage more explicitly. In our model, $H_{2}$ is used for that. The modelling is then done dynamically; for critical time periods $T$, auxiliary variables $\sigma_{j,T}$ are introduced that represent the use of CCGT power plants during periods of low renewable energy supply:
@@
At the beginning, the hydrogen storage is at $s^{H_{2}}_{0}$. In every hour after that, depending on the net energy balance, either more energy is stored or energy is taken from storage via $H_{2}$ to electricity conversion. At any point during a year, the hydrogen stored so far must be sufficient to at least cover all hydrogen demands during past/present critical periods. This can again be encoded as a MIP constraint:
@@
Case study of a German energy system model
The algorithm and all modifications were implemented and evaluated based on a fully renewable energy system model for Germany in $2045$ that includes all technologies listed in Table 4.
Two implementations were evaluated. First, a benchmark model was implemented in gurobipy that allows for comparison of all algorithms. A description of this model can be found in Appendix A Then, a more detailed $38$ node model was implemented in the open source Framework for Integrated Energy System Assessment (FINE)  41 within the ETHOS modelling suite 42 to ensure validity of the results. This includes time–series aggregation via the TSAM package 37. Since FINE currently does allow interacting with the solver during the solution process, which is necessary to add cutting planes, this was only done for MOD 2 / Synthetic time–series.
@@
$40$ years of wind and PV data in hourly resolution were taken from renewables.ninja, using data originally provided by Staffell and Pfenninger 43, 44. They provide decades of hourly open data for both wind and PV allowing for easy integration into any kind of model. The regional maximum capacity potentials for wind and PV were taken from Risch et al. 45, 46, as this is highly detailed and validated data for Germany that if freely available. The basis for geodata is the Nomenclature des Unités territoriales statistiques –- NUTS, a classification of the European Union. Level $2$ of this classification 47 is used in this work. Electricity demand data in hourly resolution for the year $2050$ for tertiary, household, transport and industry sectors were taken from the Forschungsstelle für Energiewirtschaft e. V. (FfE) 27, 28, 29, 30. On one hand, this data was selected because it covers a broad range of future electricity demands including cooling and heating demands. On the other hand, the data was calculated using the weather year $2012$ as a basis. Among the years considered in that study, 2012 was a year with an average amount of heating degree days, which fares well with our assumptions that storage levels at the first and last time steps need to be equal. The rare and extreme cold period with temperatures of $30^{\circ}$C occurring in early February of $2012$ leads to a spike in heating demand making it ideal as a basis for highly robust energy system models. In combination with the dark lull in the weather data from $1994$, this represents a severe, although short (<4<4 days) and extremely rare event in Germany and a high level of robustness for the results is to be expected. According to the German weather service 48, the lowest temperature since the beginning of recordings is $37.8^{\circ}$C. The utilised data and the developed code for either of the following calculations are publicly available via GitHub 34.
@@
Table 5: CAESAR computing cluster specifications of utilised nodes for optimising the energy system model for Germany.
@@
The full code, model and data are publicly available via GitHub 34. Any additional information required to recreate the results reported in this paper is available from the lead contact upon request.
@@ Appendix A Reference Linear Program
The following models gives a high–level explanation of the energy system model we consider. For a detailed model, we refer to the implementation provided via GitHub 34. Consider a single node model. Let $t\in T$ denote time and $p\in P=P^{prod}\cup P^{stor}\cup P^{conv}$ denote different technologies, i.e. production, conversion and storage. Then, define the following variables:
@@
Second, hydrogen storage an conversion is modelled explicitly:
@@
Appendix B Comparison of Individual Years for gurobipy Model
Figure 11 shows the results of optimising an energy system model based on multiple reference baselines.
Figure 11: Total annual cost comparison on single node model using gurobipy, no time aggregation.
Appendix C Load Shedding for Feasibility Testing
Figure 12 shows the amount of load shedding when testing the feasibility of the selected years for the other $39$ years. Two trends can be observed dependent on reference year and backup capacity built. On the one hand, when the feasibility is tested for the years $1987$ and $2010$, load shedding occurs mostly in the first half of February with just one exception in December. This is again due to the fact that the demand data used here spikes in early February caused by the low temperatures in that time in the reference year of the electricity demand. The year $2010$ only has additional load shedding in the year $1987$. The energy systems optimised for $1987$ and $2010$ are characterised by the highest amounts of PV and wind capacities installed while having comparatively low backup capacities. Both years have increased cost due to low full load hours for both wind and PV over the entire year, but periods of low electricity generation do not coincide with periods with the highest electricity demand. On the other hand, the energy system of the reference years $2009$ has increased backup capacities, therefore suffering less during cold darklull periods, but does not have enough PV and wind capacities to produce enough hydrogen and therefore has load shedding more evenly distributed over the months from October to March. The other two reference years $1996$ and $2013$ are in the middle of these two cases, suffering from both cold darklulls as well as insufficient hydrogen production, but to a lesser degree.
Figure 12: Load shedding when testing the feasibility of the $5$ selected reference years in FINE. Only the years and the months from October to March are included, where load shedding occurs.
Appendix D Techno–Economic parameters
@@
Appendix E Description of Modifications
@@
Appendix F Evaluation of Ineffective Modifications
None of the modifications that are outlined in Figure E are able to generate robust solutions on their own. Reasons for that are outlined below:

### v4

@@ A Novel Algorithm for Computing Robust ESMs
If an ESM displays supply gaps during feasibility testing, its CAPEX optimisation problem is modified to obtain a robust solution. Consider a solution $\delta$ to $COMP^{i}_{j}$. In this solution, $1^{\intercal}\delta_{i}^{x}$ units of energy are missing. This information needs to be integrated into the original CAPEX problem to enable the ESM to provide the extra energy supply required to meet demand.
Multiple modifications were iteratively developed to capture different sources of uncertainty. Initial testing showed a subset of them to be ineffective. In order to also report null–results, these modifications are outlined in Appendix C. The three promising modifications are inspired by three physical phenomena relevant to energy systems: 1) The importance of peak demands for energy system design 56, 32, 2) The existence of extended critical time periods 48, 46, 21, 22 and 3) variations in total yearly energy supply and demand across different years 8.
For MOD 1 / Demand increase, update the demand vector $b_{j}$ of the starting year $j$ by adding the demand-supply gap $\delta$ to it, i.e.
@@
This adds the missing energy supply exactly when it is needed. One observation during initial testing was that this can lead to very large peaks in artificial demand, especially if done repeatedly over multiple iterations, leading to excessive battery installation. To counteract this, the artificial, additional demand can be divided up between neighbouring time periods, i.e. update $b$ and smooth $\delta$, i.e. $b^{\prime}=b+f(\delta)$ for some function $f:\mathbb{R}^{n}_{\geq 0}\mapsto\mathbb{R}^{n}_{\geq 0}\setminus\emptyset$ that changes $b$ locally. Figure 12 in Appendix C illustrates the effects of smoothing. Throughout our work, we differentiate between smoothed and non–smoothed MOD 1 / Demand increase.
Since feasibility tests show that supply gaps usually occur in adjacent time steps, those are clustered to identify time periods with potential supply gaps, i.e. critical time periods. Critical time periods are usually characterised by a combination of low electricity generation from renewables together with increased electricity demand, so–called (cold) dark lulls. A hierarchical clustering was chosen, as this has been shown to be effective for time–series aggregation previously 28, 1, and it allows for easy visual inspection and clustering along adjacent time–steps 26, 40. For each time period, check whether the optimised capacities provide sufficient energy. For this, calculate the potential electricity generation of the selected ESM and compare this with the total demand in the time period. [^4] Note that a positive average supply gap means that the investigated energy systems lack dispatchable capacity, and thus the feasibility testing of the respective time periods will always find a supply gap. However, a time period with negative average hourly supply gaps can still contain supply gaps that are only found during feasibility testing. This happens if either total annual electricity generation is too low to cover overall demand, or if there are some demand peaks that can not be covered. For calculating the potential electricity generation, full load is assumed for generation and backup capacities. Let $b_{t,j}$ be the demand at hour $t\in T$, and $a_{t,j}^{k}\cdot x_{s}$ the time–series data and the calculated capacities of the energy system component $p\in P$ in year $u_{j}$, respectively. $P_{sup,con}=P_{sup}\cup P_{con}$ is used to denote that summation only happens over supply and conversion technologies. Then, for time interval $T\subseteq\mathcal{T}$, the average hourly supply gap $\Delta(T)_{j}$ for some year $u_{j}^{\prime}$ is given via
@@
for some critical time period $T$ and years $j,j^{\prime}$.
@@
If this does not suffice to achieve feasibility for some year $j$, we model long-term energy storage more explicitly. In our model, $H_{2}$ is used for that. The modelling is then done dynamically; for critical time periods $T$, auxiliary variables $\sigma_{j,T}$ are introduced that represent the use of CCGT power plants during periods of low renewable energy supply:
@@
At the beginning, the hydrogen storage is at $s^{H_{2}}_{0}$. In every hour after that, depending on the net energy balance, either more energy is stored or energy is taken from storage via $H_{2}$ to electricity conversion. At any point during a year, the hydrogen stored so far must be sufficient to at least cover all hydrogen demands during past/present critical periods. This can again be encoded as a MILP constraint:
@@
Case study of a German ESM
The algorithm and all modifications were implemented and evaluated based on a fully renewable ESM for Germany in $2045$ that includes all technologies listed in Table 4.
Two implementations were evaluated. First, a benchmark model was implemented in gurobipy that allows for comparison of all algorithms. A description of this model can be found in Appendix A Then, a more detailed $38$ node model was implemented in the open source Framework for Integrated Energy System Assessment (ETHOS.FINE)  57 within the ETHOS modelling suite 23 to ensure validity of the results. This includes time–series aggregation via the TSAM package 28. Since ETHOS.FINE currently does allow interacting with the solver during the solution process, which is necessary to add cutting planes, this was only done for MOD 2 / Synthetic time–series.
@@
$40$ years of wind and PV data in hourly resolution were taken from renewables.ninja, using data originally provided by Staffell and Pfenninger 38, 51. They provide decades of hourly open data for both wind and PV allowing for easy integration into any kind of model. The regional maximum capacity potentials for wind and PV were taken from Risch et al. 43, 42, as this is highly detailed and validated data for Germany that if freely available. The basis for geodata is the Nomenclature des Unités territoriales statistiques –- NUTS, a classification of the European Union. Level $2$ of this classification 9 is used in this work. Electricity demand data in hourly resolution for the year $2050$ for tertiary, household, transport and industry sectors were taken from the Forschungsstelle für Energiewirtschaft e. V. (FfE) 17, 15, 18, 16. On one hand, this data was selected because it covers a broad range of future electricity demands including cooling and heating demands. On the other hand, the data was calculated using the weather year $2012$ as a basis. Among the years considered in that study, 2012 was a year with an average amount of heating degree days, which fares well with our assumptions that storage levels at the first and last time steps need to be equal. The rare and extreme cold period with temperatures of $30^{\circ}$C occurring in early February of $2012$ leads to a spike in heating demand making it ideal as a basis for highly robust ESMs. In combination with the dark lull in the weather data from $1994$, this represents a severe, although short (<4<4 days) and extremely rare event in Germany and a high level of robustness for the results is to be expected. According to the German weather service 11, the lowest temperature since the beginning of recordings is $37.8^{\circ}$C. The utilised data and the developed code for either of the following calculations are publicly available via GitHub 33. The cluster specifications are outlined in Table 5.
@@
Table 5: CAESAR computing cluster specifications of utilised nodes for optimising the ESM for Germany.
QUANTIFICATION AND STATISTICAL ANALYSIS
There are no quantification or statistical analyses to include in this study.
@@
Materials availability
This study did not generate new materials.
@@
The full code, model and data are publicly available via GitHub 33. Any additional information required to recreate the results reported in this paper is available from the lead contact upon request.
@@ Appendix A Reference Linear Program
The following models gives a high–level explanation of the ESM we consider. For a detailed model, we refer to the implementation provided via GitHub 33. Consider a single node model. Let $t\in T$ denote time and $p\in P=P^{prod}\cup P^{stor}\cup P^{conv}$ denote different technologies, i.e. production, conversion and storage. Then, define the following variables:
@@
Second, hydrogen storage and conversion is modelled explicitly:
@@
Appendix B Techno–Economic parameters
@@
Additionally, in ETHOS.FINE we also assumed a self discharge of $0.004\%$ per hour for Li–ion batteries.
Appendix C Description of Modifications
@@
Appendix D Evaluation of Ineffective Modifications
None of the modifications that are outlined in Section C are able to generate robust solutions on their own. Reasons for that are outlined below: